OpenStudy (anonymous):

Find the geometric mean between 5 and 8?

OpenStudy (amistre64):

geometric means simple enough. for 2 numbers, the geoMean is the square root of their product

OpenStudy (amistre64):

\[geoMean(a,b)=\sqrt{ab}\]

OpenStudy (anonymous):

6.3?

OpenStudy (amistre64):

are they asking you to round it?

OpenStudy (anonymous):

nope

OpenStudy (amistre64):

\[\sqrt{5*4*2}\] \[\sqrt{4*10}\] \[\sqrt{4}*\sqrt{10}\] \[2\sqrt{10}\]

OpenStudy (anonymous):

so im wrong?

OpenStudy (amistre64):

6.3 is a rational value; 2sqrt(10) is an irrational value since they do not equate to each other, and you are not asked to round anything; then your 6.3 is wrong

OpenStudy (anonymous):

then how whats the right anwer, or how do i solve it?